OpenStudy (anonymous):

A quadratic function is given. f(x) = −x^2 − 4x + 4 (a) Express the quadratic function in standard form. f(x) = ?

OpenStudy (anonymous):

This needs to be in form to find the vertex. But I'm lost on how to do this.

OpenStudy (anonymous):

for finding vertex, standard form is not necessary. @hartnn am i correct?

hartnn (hartnn):

since it need to be in the form of vertex, it should be like f(x)=a(x-h)^2+k then vertex is (h,k)

hartnn (hartnn):

so complete the square....

OpenStudy (anonymous):

Yeah. I just don't know how to convert it from this form to f(x) = a(x-h)^2+k

hartnn (hartnn):

factor out -1, u get -1(x^2+4x-4) now half the middle term like divide the co-efficient of x by 2 what u get ?

OpenStudy (anonymous):

or can we use direct formula of vertex (-b/2a , -D/4a)

OpenStudy (anonymous):

4/2 = 2 2^2=4 (square it, correct?)

hartnn (hartnn):

yes, now add and subtract 4 in that function

OpenStudy (anonymous):

-1(x^2+4x -4 +4) -4 ?

hartnn (hartnn):

umm, almost... -1((x^2+4x +4) -4 -4) got it ?

hartnn (hartnn):

i just rearranged the terms and put the brackets

OpenStudy (anonymous):

Yes! Makes sense. So would it be: -1(x+2)^2-8 ?

hartnn (hartnn):

be careful with signs -1((x+2)^2-8 ) -1(x+2)^2+8

OpenStudy (anonymous):

How are you getting positive 8? are you distributing?

hartnn (hartnn):

yes, i need 'k' separate so i distributed -1

OpenStudy (anonymous):

Oh! ok. Thank you very much!

hartnn (hartnn):

welcome ^_^